OpenStudy (anonymous):

Factor this expression as a difference of squares: 4-(2a+3b)^2

OpenStudy (anonymous):

\(x^2-y^2=(x+y)(x-y)\) make a direct substitution with \(x=2, y=2a+3b\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^2 - b^2 = (a + b) (a - b)}\) \( \color{Black}{\Rightarrow 2^2 - (2a + 3b)^2 = ....?}\)

OpenStudy (anonymous):

You guys are ALWAYS the first to answer.

OpenStudy (anonymous):

Im terrible at math.. Please help me... :(

OpenStudy (saranya):

4 − ( 2a + 3b )² Write 4 in the under rooted form = ( 2 )² − ( 2a + 3b )² Factorise as the difference of 2 squares = a² − b² = ( a + b )( a − b ) = [ 2 + ( 2a + 3b ) ][ 2 − ( 2a + 3b ) ] = ( 2 + 2a + 3b )( 2 - 2a - 3b ) Arrange = ( 2a + 3b + 2 )( 2 − 2a − 3b )

OpenStudy (anonymous):

so the answer is...??

OpenStudy (anonymous):

(2a + 3b + 2)(2 - 2a - 3b) Oh my, you are terrible at math.